OpenStudy (anonymous):

x ÷ 0 = 0 always never sometimes

OpenStudy (ipwnbunnies):

Never. Anything divided by zero is undefined/indeterminate.

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

You can check this by trying it in your calculator if you like, it will say error :P

OpenStudy (whpalmer4):

Think about it: a/b = c implies that c*b = a Now, if we divide 1/0, if we say that equals 0, then that implies that 0*0=1! But that doesn't make sense, because clearly 0*0=0. 0 times any number equals 0.